# Large Language Models for Information Retrieval: A Survey

### v3

@@ Abstract
Large Language Models; Information Retrieval; Query Rewrite; Rerank; Reader; Fine-tuning; Prompting
@@ 1 Introduction
The core function of an IR system is retrieval, which aims to determine the relevance between a user-issued query and the content to be retrieved, including various types of information such as texts, images, music, and more. For the scope of this survey, we concentrate solely on reviewing those text retrieval systems, in which query-document relevance is commonly measured by their matching score. [^3] The term “document” will henceforth refer to any text-based content subject to retrieve, including both long articles and short passages. Given that IR systems operate on extensive repositories, the efficiency of retrieval algorithms becomes of paramount importance. To improve the user experience, the retrieval performance is enhanced from both the upstream (query reformulation) and downstream (reranking and reading) perspectives. As an upstream technique, query reformulation is designed to refine user queries so that they are more effective at retrieving relevant documents [10, 11]. With the recent surge in the popularity of conversational search, this technique has received increasing attention. On the downstream side, reranking approaches are developed to further adjust the document ranking [12, 13, 14]. In contrast to the retrieval stage, reranking is performed only on a limited set of relevant documents, already retrieved by the retriever. Under this circumstance, the emphasis is placed on achieving higher performance rather than keeping higher efficiency, allowing for the application of more complex approaches in the reranking process. Additionally, reranking can accommodate other specific requirements, such as personalization [15, 16, 17, 18] and diversification [19, 20, 21, 22]. Following the retrieval and reranking stages, a reading component is incorporated to summarize the retrieved documents and deliver a concise document to users [23, 24]. While traditional IR systems typically require users to gather and organize relevant information themselves; however, the reading component is an integral part of new IR systems such as New Bing, [^4] New Bing, https://www.bing.com/new streamlining users’ browsing experience and saving valuable time.
@@
Large language models (LLMs) have recently emerged as transformative forces across various research fields, such as natural language processing (NLP) [33, 34, 35], recommender systems [36, 37, 38, 39], finance [40], and even molecule discovery [41]. These cutting-edge LLMs are primarily based on the Transformer architecture and undergo extensive pre-training on diverse textual sources, including web pages, research articles, books, and codes. As their scale continues to expand (including both model size and data volume), LLMs have demonstrated remarkable advances in their capabilities. On the one hand, LLMs have exhibited unprecedented proficiency in language understanding and generation, resulting in responses that are more human-like and better align with human intentions. On the other hand, the larger LLMs have shown impressive emergent abilities when dealing with complex tasks [42], such as generalization and reasoning skills. Notably, LLMs can effectively apply their learned knowledge and reasoning abilities to tackle new tasks with just a few task-specific demonstrations or appropriate instructions [43, 44]. Furthermore, advanced techniques, such as in-context learning, have significantly enhanced the generalization performance of LLMs without requiring fine-tuning on specific downstream tasks [34]. This breakthrough is particularly valuable, as it reduces the need for extensive fine-tuning while attaining remarkable task performance. Powered by prompting strategies such as chain-of-thought, LLMs can generate outputs with step-by-step reasoning, navigating complex decision-making processes [45]. Leveraging the impressive power of LLMs can undoubtedly improve the performance of IR systems. By incorporating these sophisticated language models, IR systems can provide users with more accurate responses, ultimately reshaping the landscape of information access and retrieval.
@@
Our survey provides an insightful exploration of the intersection between LLMs and IR systems, covering key perspectives such as query rewriters, retrievers, rerankers, and readers (as shown in Figure 1). [^5] As yet, there has not been a formal definition for LLMs. In this paper, we mainly focus on models with more than 1B parameters. We also notice that some methods do not rely on such strictly defined LLMs, but due to their representativeness, we still include an introduction to them in this survey. We also include some recent studies that leverage LLMs as search agents to perform various IR tasks. This analysis enhances our understanding of LLMs’ potential and limitations in advancing the IR field. For this survey, we create a Github repository by collecting the relevant papers and resources about LLM4IR. [^6] https://github.com/RUC-NLPIR/LLM4IR-Survey We will continue to update the repository with newer papers. This survey will also be periodically updated according to the development of this area. We notice that there are several surveys for PLMs, LLMs, and their applications (e.g., AIGC or recommender systems) [46, 47, 48, 49, 50, 51, 52]. Among these, we highly recommend the survey of LLMs [52], which provides a systematic and comprehensive reference to many important aspects of LLMs. Compared with them, we focus on the techniques and methods for developing and applying LLMs for IR systems. In addition, we notice a perspective paper discussing the opportunity of IR when meeting LLMs [53]. It would be an excellent supplement to this survey regarding future directions.
The remaining part of this survey is organized as follows: Section 2 introduces the background for IR and LLMs. Section 345VII respectively review recent progress from the four perspectives of query rewriter, retriever, reranker, and reader, which are four key components of an IR system. Then, Section 8 discusses some potential directions in future research. Finally, we conclude the survey in Section 9 by summarizing the major findings.
@@ 2.1 Information Retrieval
Information retrieval (IR), as an essential branch of computer science, aims to efficiently retrieve information relevant to user queries from a large repository. Generally, users interact with the system by submitting their queries in textual form. Subsequently, IR systems undertake the task of matching and ranking these user-supplied queries against an indexed database, thereby facilitating the retrieval of the most pertinent results.
The field of IR has witnessed significant advancement with the emergence of various models over time. One such early model is the Boolean model, which employs Boolean logic operators to combine query terms and retrieve documents that satisfy specific conditions [25]. Based on the “bag-of-words” assumption, the vector space model [26] represents documents and queries as vectors in term-based space. Relevance estimation is then performed by assessing the lexical similarity between the query and document vectors. The efficiency of this model is further improved through the effective organization of text content using the inverted index. Moving towards more sophisticated approaches, statistical language models have been introduced to estimate the likelihood of term occurrences and incorporate context information, leading to more accurate and context-aware retrieval [27, 54]. In recent years, the neural IR [30, 55, 56] paradigm has gained considerable attention in the research community. By harnessing the powerful representation capabilities of neural networks, this paradigm can capture semantic relationships between queries and documents, thereby significantly enhancing retrieval performance.
@@
Query Rewriter is an essential IR module that seeks to improve the precision and expressiveness of user queries. Positioned at the early stage of the IR pipeline, this module assumes the crucial role of refining or modifying the initial query to align more accurately with the user’s information requirements. As an integral part of query rewriting, query expansion techniques, with pseudo relevance feedback being a prominent example, represent the mainstream approach to achieving query expression refinement. In addition to its utility in improving search effectiveness across general scenarios, the query rewriter finds application in diverse specialized retrieval contexts, such as personalized search and conversational search, thus further demonstrating its significance.
@@ 2.2 Large Language Models
Language models (LMs) are designed to calculate the generative likelihood of word sequences by taking into account the contextual information from preceding words, thereby predicting the probability of subsequent words. Consequently, by employing certain word selection strategies (such as greedy decoding or random sampling), LMs can proficiently generate natural language texts. Although the primary objective of LMs lies in text generation, recent studies [57] have revealed that a wide array of natural language processing problems can be effectively reformulated into a text-to-text format, thus rendering them amenable to resolution through text generation. This has led to LMs becoming the de facto solution for the majority of text-related problems.
The evolution of LMs can be categorized into four primary stages, as discussed in prior literature [52]. Initially, LMs were rooted in statistical learning techniques and were termed statistical language models. These models tackled the issue of word prediction by employing the Markov assumption to predict the subsequent word based on preceding words. Thereafter, neural networks, particularly recurrent neural networks (RNNs), were introduced to calculate the likelihood of text sequences and establish neural language models. These advancements made it feasible to utilize LMs for representation learning beyond mere word sequence modeling. ELMo [58] first proposed to learn contextualized word representations through pre-training a bidirectional LSTM (biLSTM) network on large-scale corpora, followed by fine-tuning on specific downstream tasks. Similarly, BERT [59] proposed to pre-train a Transformer [60] encoder with a specially designed Masked Language Modeling (MLM) task and Next Sentence Prediction (NSP) task on large corpora. These studies initiated a new era of pre-trained language models (PLMs), with the “pre-training then fine-tuning” paradigm emerging as the prevailing learning approach. Along this line, numerous generative PLMs (e.g., GPT-2 [33], BART [61], and T5 [57]) have been developed for text generation problems including summarization, machine translation, and dialogue generation. Recently, researchers have observed that increasing the scale of PLMs (e.g., model size or data amount) can consistently improve their performance on downstream tasks (a phenomenon commonly referred to as the scaling law [62, 63]). Moreover, large-sized PLMs exhibit promising abilities (termed emergent abilities [42]) in addressing complex tasks, which are not evident in their smaller counterparts. Therefore, the research community refers to these large-sized PLMs as large language models (LLMs).
Figure 2: The evolution of LLMs (encoder-decoder and decoder-only structures).
As shown in Figure 2, existing LLMs can be categorized into two groups based on their architectures: encoder-decoder [61, 57, 64, 65, 66, 67, 68, 69] and decoder-only [33, 34, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 35, 80] models. The encoder-decoder models incorporate an encoder component to transform the input text into vectors, which are then employed for producing output texts. For example, T5 [57] is an encoder-decoder model that converts each natural language processing problem into a text-to-text form and resolves it as a text generation problem. In contrast, decoder-only models, typified by GPT, rely on the Transformer decoder architecture. It uses a self-attention mechanism with a diagonal attention mask to generate a sequence of words from left to right. Building upon the success of GPT-3 [34], which is the first model to encompass over 100B parameters, several noteworthy models have been inspired, including GPT-J, BLOOM [78], OPT [75], Chinchilla [81], and LLaMA [35]. These models follow the similar Transformer decoder structure as GPT-3 and are trained on various combinations of datasets.
@@
Figure 3: An example of LLM-based query rewriting for ad-hoc search. The example is cited from the Query2Doc paper [86]. LLMs are used to generate a passage to supplement the original query, where $N=0$ and $N>0$ correspond to zero-shot and few-shot scenarios.
@@ 3 Query Rewriter
Query rewriting in modern IR systems is essential for improving search query effectiveness and accuracy. It reformulates users’ original queries to better match search results, alleviating issues like vague queries or vocabulary mismatches between the query and target documents. This task goes beyond mere synonym replacement, requiring an understanding of user intent and query context, particularly in complex searches like conversational queries. Effective query rewriting enhances search engine performance.
Traditional methods for query rewriting improve retrieval performance by expanding the initial query with information from highly-ranked relevant documents. Mainly-used methods include relevance feedback [87, 88, 89, 90, 91, 92], word-embedding based methods [93, 94] etc. However, the limited ability of semantic understanding and comprehension of user search intent limits their performance in capturing the full scope of user intent.
Recent advancements in LLMs present promising opportunities to boost query rewriting capabilities. On one hand, given the context and subtleties of a query, LLMs can provide more accurate and contextually relevant rewrites. On the other hand, LLMs can leverage their extensive knowledge to generate synonyms and related concepts, enhancing queries to cover a broader range of relevant documents, thereby effectively addressing the vocabulary mismatch problem. In the following sections, we will introduce the recent works that employ LLMs in query rewriting.
@@
Figure 4: An example of LLM-based query rewriting for conversational search. The example is cited from LLMCS [95]. The LLM is used to generate a query based on the demonstrations and previous search context. Additional responses are required to be generated for improving the query understanding. $N=0$ and $N>0$ correspond to zero-shot and few-shot scenarios.
3.1 Rewriting Scenario
Query rewriting typically serves two scenarios: ad-hoc retrieval, which mainly addresses vocabulary mismatches between queries and candidate documents, and conversational search, which refines queries based on evolving conversations. The upcoming section will delve into the role of query rewriting in these two domains and explore how LLMs enhance this process.
3.1.1 Ad-hoc Retrieval
In ad-hoc retrieval, queries are often short and ambiguous. In such scenarios, the main objectives of query rewriting include adding synonyms or related terms to address vocabulary mismatches and clarifying ambiguous queries to more accurately align with user intent. From this perspective, LLMs have inherent advantages in query rewriting. Primarily, LLMs have a deep understanding of language semantics, allowing them to capture the meaning of queries more effectively. Besides, LLMs can leverage their extensive training on diverse datasets to generate contextually relevant synonyms and expand queries, ensuring broader and more precise search result coverage. Additionally, studies have shown that LLMs’ integration of external factual corpora  [96, 97, 98, 99] and thoughtful model design [100] further enhance their accuracy in generating effective query rewrites, especially for specific tasks.
Currently, there are many studies leveraging LLMs to rewrite queries in adhoc retrieval. We introduce the typical method Query2Doc [86] as an example. As shown in Figure 3, Query2Doc prompts the LLMs to generate a relevant passage according to the original query (“when was pokemon green released?”). Subsequently, the original query is expanded by incorporating the generated passage. The retriever module uses this new query to retrieve a list of relevant documents. Notably, the generated passage contains additional detailed information, such as “Pokemon Green was released in Japan on February 27th”, which effectively mitigates the “vocabulary mismatch” issue to some extent.
In addition to addressing the “vocabulary mismatch” problem [101, 102, 96, 97, 98, 99], other works utilize LLMs for different challenges in ad-hoc retrieval. For instance, PromptCase [103] leverages LLMs in legal case retrieval to simplify complex queries into more searchable forms. This involves using LLMs to identify legal facts and issues, followed by a prompt-based encoding scheme for effective language model encoding.
3.1.2 Conversational Search
Query rewrites in conversational search play a pivotal role in enhancing the search experience. Unlike traditional queries in ad-hoc retrieval, conversational search involves a dialogue-like interaction, where the context and user intent evolve with each interaction. In conversational search, query rewriting involves understanding the entire conversation’s context, clarifying any ambiguities, and personalizing responses based on user history. The process includes dynamic query expansion and refinement based on dialogue information. This makes conversational query rewriting a sophisticated task that goes beyond traditional search, focusing on natural language understanding and user-centric interaction.
In the era of LLMs, leveraging LLMs in conversational search tasks offers several advantages. First, LLMs possess strong contextual understanding capabilities, enabling them to better comprehend users’ search intent within the context of multi-turn conversations between users and the system. Second, LLMs exhibit powerful generation abilities, allowing them to simulate dialogues between users and the system, thereby facilitating more robust search intent modeling.
The LLMCS framework [95] is a pioneering approach that employs LLMs to effectively extract and understand user search intent within conversational contexts. As illustrated in their work, LLMCS uses LLMs to produce both query rewrites and extensive hypothetical system responses from various perspectives. These outputs are combined into a comprehensive representation that effectively captures the user’s full search intent. The experimental results show that including detailed hypothetical responses with concise query rewrites markedly improves search performance by adding more plausible search intent. Ye et al. [104] claims that human query rewrite may lack sufficient information for optimal retrieval performance. It defines four essential properties for well-formed LLM-generated query rewrites. Results show that their method informative query rewrites can yield substantially improved retrieval performance compared to human rewrites.
Besides, LLMs can be used as a data expansion tool in conversational dense retrieval. Attributed to the high cost of producing hand-written dialogues, data scarcity presents a significant challenge in the domain of conversational search. To address this problem, CONVERSER [105] employs LLMs to generate synthetic passage-dialogue pairs through few-shot demonstrations. Furthermore, it efficiently trains a dense retriever using a minimal dataset of six in-domain dialogues, thus mitigating the issue of data sparsity.
3.2 Rewriting Knowledge
Query rewriting typically necessitates additional corpora for refining initial queries. Considering that LLMs incorporate world knowledge in their parameters, they are naturally capable of rewriting queries. We refer to these methods, which rely exclusively on the intrinsic knowledge of LLMs, as LLM-only methods. While LLMs encompass a broad spectrum of knowledge, they may be inadequate in specialized areas. Furthermore, LLMs can introduce concept drift, leading to noisy relevance signals. To address this issue, some methods incorporate domain-specific corpora to provide more detailed and relevant information in query rewriting. We refer to methods enhanced by domain-specific corpora to boost LLM performance as corpus-enhanced LLM-based methods. In this section, we will introduce these two methods in detail.
3.2.1 LLM-only methods
LLMs are capable of storing knowledge within their parameters, making it a natural choice to capitalize on this knowledge for the purpose of query rewriting. As a pioneering work in LLM-based query rewriting, HyDE [101] generates a hypothetical document by LLMs according to the given query and then uses a dense retriever to retrieve relevant documents from the corpus that are relevant to the generated document. Query2doc [86] generates pseudo documents via prompting LLMs with few-shot demonstrations, and then expands the query with the generated pseudo document. Furthermore, the influence of different prompting methods and various model sizes on query rewriting has also been investigated [102]. To better accommodate the frozen retriever and the LLM-based reader, a small language model is employed as the rewriter that is trained using reinforcement learning techniques with the rewards provided by the LLM-based reader [100]. GFF [106] presents a “Generate, Filter, and Fuse” method for query expansion. It employs an LLM to create a set of related keywords via a reasoning chain. Then, a self-consistency filter is used to identify the most important keywords, which are concatenated with the original queries for the downstream reranking task.
It is worth noting that though the designs of these methods are different, all of them rely on the world knowledge stored in LLMs without additional corpora.
3.2.2 Corpus-enhanced LLM-based methods
Although LLMs exhibit remarkable capabilities, the lack of domain-specific knowledge may lead to the generation of hallucinatory or irrelevant queries. To address this issue, recent studies [96, 97, 98, 99] have proposed a hybrid approach that enhances LLM-based query rewriting methods with an external document corpus.
Why incorporate a document corpus? The integration of a document corpus offers several notable advantages. Firstly, it boosts relevance by using relevant documents to refine query generation, reducing irrelevant content and improving contextually appropriate outputs. Second, enhancing LLMs with up-to-date information and specialized knowledge in specific fields enables them to effectively deal with queries that are both current and specific to certain domains.
How to incorporate a document corpus? Thanks to the flexibility of LLMs, various paradigms have been proposed to incorporate a document corpus into LLM-based query rewriting, which can be summarized as follows.
$\bullet$ Late fusion of LLM-based re-writing and pseudo relevance feedback (PRF) retrieval results. Traditional PRF methods leverage relevant documents retrieved from a document corpus to rewrite queries, which restricts the query to the information contained in the target corpus. On the contrary, LLM-based rewriting methods provide external context not present in the corpus, which is more diverse. Both approaches have the potential to independently enhance retrieval performance. Therefore, a straightforward strategy for combining them is using a weighted fusion method for retrieval results [99].
$\bullet$ Combining retrieved relevant documents in the prompts of LLMs. In the era of LLMs, incorporating instructions within the prompts is the most flexible method for achieving specific functionalities. QUILL [97] and CAR [107] illustrate how retrieval augmentation of queries can provide LLMs with context that significantly enhances query understanding. LameR [108] takes this further by using LLM expansion to improve the simple BM25 retriever, introducing a retrieve-rewrite-retrieve framework. Experimental results reveal that even basic term-based retrievers can achieve comparable performance when paired with LLM-based rewriters. Additionally, InteR [98] proposes a multi-turn interaction framework between search engines and LLMs. This enables search engines to expand queries using LLM-generated insights, while LLMs refine prompts using relevant documents sourced from the search engines.
$\bullet$ Enhancing factuality of generative relevance feedback (GRF) by pseudo relevance feedback (PRF). Although generative documents are often relevant and diverse, they exhibit hallucinatory characteristics. In contrast, traditional documents are generally regarded as reliable sources of factual information. Motivated by this observation, GRM [96] proposes a novel technique known as relevance-aware sample estimation (RASE). RASE leverages relevant documents retrieved from the collection to assign weights to generated documents. In this way, GRM ensures that relevance feedback is not only diverse but also maintains a high degree of factuality.
3.3 Rewriting Approaches
There are three main approaches used for leveraging LLMs in query rewriting: prompting methods, fine-tuning, and knowledge distillation. Prompting methods involve using specific prompts to direct LLM output, providing flexibility and interpretability. Fine-tuning adjusts pre-trained LLMs on specific datasets or tasks to improve domain-specific performance, mitigating the general nature of LLM world knowledge. Knowledge distillation, on the other hand, transfers LLM knowledge to lightweight models, simplifying the complexity associated with retrieval augmentation. In the following section, we will introduce these three methods in detail.

### v4

@@ Abstract
Large Language Models; Information Retrieval; Query Rewriter; Reranking; Reader; Fine-tuning; Prompting
@@ 1 Introduction
The core function of an IR system is retrieval, which aims to determine the relevance between a user-issued query and the content to be retrieved, including various types of information such as texts, images, music, and more. For the scope of this survey, we concentrate solely on reviewing those text retrieval systems, in which query-document relevance is commonly measured by their matching score. [^3] The term “document” will henceforth refer to any text-based content subject to retrieve, including both long articles and short passages. Given that IR systems operate on extensive repositories, the efficiency of retrieval algorithms becomes of paramount importance. To improve the user experience, the retrieval performance is enhanced from both the upstream (query reformulation) and downstream (reranking and reading) perspectives. As an upstream technique, query reformulation is designed to refine user queries so that they are more effective at retrieving relevant documents [10, 11]. With the recent surge in the popularity of conversational search, this technique has received increasing attention. On the downstream side, reranking approaches are developed to further adjust the document ranking [12, 13, 14]. In contrast to the retrieval stage, reranking is performed only on a limited set of relevant documents, already retrieved by the retriever. Under this circumstance, the emphasis is placed on achieving higher performance rather than keeping higher efficiency, allowing for the application of more complex approaches in the reranking process. Additionally, reranking can accommodate other specific requirements, such as personalization [15, 16, 17, 18] and diversification [19, 20, 21, 22]. Following the retrieval and reranking stages, a reading component is incorporated to summarize the retrieved documents and deliver a concise document to users [23, 24]. While traditional IR systems typically require users to gather and organize relevant information themselves; however, the reading component is an integral part of new IR systems such as New Bing, [^4] New Bing, https://www.bing.com/new improving users’ browsing experience and saving valuable time.
@@
Large language models (LLMs) have recently emerged as transformative forces across various research fields, such as natural language processing (NLP) [33, 34, 35], recommender systems [36, 37, 38, 39], finance [40], and even molecule discovery [41]. These cutting-edge LLMs are primarily based on the Transformer architecture and undergo extensive pre-training on diverse textual sources, including web pages, research articles, books, and codes. As their scale continues to expand (including both model size and data volume), LLMs have demonstrated remarkable advances in their capabilities. On the one hand, LLMs have exhibited unprecedented proficiency in language understanding and generation, resulting in responses that are more human-like and better aligned with human intentions. On the other hand, the larger LLMs have shown impressive emergent abilities when dealing with complex tasks [42], such as generalization and reasoning skills. Leveraging the impressive power of LLMs can undoubtedly improve the performance of IR systems. By incorporating these advanced language models, IR systems can provide users with more accurate responses, ultimately reshaping the landscape of information access and retrieval.
@@
Our survey provides an insightful exploration of the intersection between LLMs and IR systems, covering key perspectives such as query rewriters, retrievers, rerankers, and readers (as shown in Figure 1). [^5] As yet, there has not been a formal definition for LLMs. In this paper, we mainly focus on models with more than 1B parameters. We also notice that some methods do not rely on such strictly defined LLMs, but due to their representativeness, we still include an introduction to them in this survey. We also include some recent studies that leverage LLMs as search agents to perform various IR tasks. This analysis enhances our understanding of LLMs’ potential and limitations in advancing the IR field. For this survey, we create a Github repository by collecting the relevant papers and resources about applying LLMs for IR tasks (LLM4IR). [^6] https://github.com/RUC-NLPIR/LLM4IR-Survey We will continue to update the repository with newer papers. This survey will also be periodically updated according to the development of this area. We notice that there are several surveys for PLMs, LLMs, and their applications (e.g., AIGC or recommender systems) [43, 44, 45, 46, 47, 48, 49]. Compared with them, we focus on the techniques and methods for developing and applying LLMs for IR systems. In addition, we suggest reading the strategy report from the Chinese IR community [50], which discusses the opportunity and future directions of IR in the era of LLMs, and we think it is an excellent supplement to this survey.
The remaining part of this survey is organized as follows: Section 2 introduces the background for IR and LLMs. Section 345VII respectively review recent progress from the four perspectives of query rewriter, retriever, reranker, and reader, which are four key components of an IR system. Section 7 introduces recent studies of search agents. Then, Section 8 discusses some potential directions in future research. Finally, we conclude the survey in Section 9 by summarizing the major findings.
@@ 2.1 Information Retrieval
Information retrieval (IR), as an essential branch of computer science, aims to efficiently retrieve information relevant to user queries from a large repository. Generally, users interact with an IR system by submitting their queries in textual form. Subsequently, IR systems undertake the task of matching and ranking these user-supplied queries against an indexed database, thereby facilitating the retrieval of the most pertinent results.
The field of IR has witnessed significant advancement with the emergence of various models over time. One such early model is the Boolean model, which employs Boolean logic operators to combine query terms and retrieve documents that satisfy specific conditions [25]. Based on the “bag-of-words” assumption, the vector space model [26] represents documents and queries as vectors in term-based space. Relevance estimation is then performed by assessing the lexical similarity between the query and document vectors. The efficiency of this model is further improved through the effective organization of text content using the inverted index. Moving towards more sophisticated approaches, statistical language models have been introduced to estimate the likelihood of term occurrences and incorporate context information, leading to more accurate and context-aware retrieval [27, 51]. In recent years, the neural IR paradigm has gained considerable attention in the research community [30, 52, 53]. By harnessing the powerful representation capabilities of neural networks, this paradigm can capture semantic relationships between queries and documents, thereby significantly enhancing retrieval performance.
@@
Query Rewriter is an essential IR module that seeks to improve the precision and expressiveness of user queries. Positioned at the early stage of the IR pipeline, this module assumes the crucial role of refining or modifying the initial query to align more accurately with the user’s information requirements. As an integral part of query rewriter, query expansion techniques, with pseudo relevance feedback being a prominent example, represent the mainstream approach to achieving query expression refinement. In addition to its utility in improving search effectiveness across general scenarios, the query rewriter finds application in diverse specialized retrieval contexts, such as personalized search and conversational search, thus further demonstrating its significance.
@@ 2.2 Large Language Models
Language models (LMs) are designed to understand or generate human language by taking into account the contextual information from word sequences. The evolution from statistical language models to neural language models makes it feasible to utilize LMs for representation learning beyond mere word sequence modeling. Peters et al. [54] first proposed to learn contextualized word representations through pre-training a bidirectional LSTM (biLSTM) network on large-scale corpora, followed by fine-tuning on specific downstream tasks. Similarly, Devlin et al. [55] proposed to pre-train a Transformer [56] encoder with a specially designed Masked Language Modeling (MLM) task and Next Sentence Prediction (NSP) task on large corpora. These studies initiated a new era of pre-trained language models (PLMs), with the “pre-training then fine-tuning” paradigm emerging as the prevailing learning approach. Along this line, numerous generative PLMs (e.g., GPT-2 [33], BART [57], and T5 [58]) have been developed for text generation problems including summarization, machine translation, and dialogue generation. Recently, researchers have observed that increasing the scale of PLMs (e.g., model size or data amount) can consistently improve their performance on downstream tasks (a phenomenon commonly referred to as the scaling law [59, 60]). Moreover, large-sized PLMs exhibit promising abilities (termed emergent abilities [42]) in addressing complex tasks, which are not evident in their smaller counterparts. Therefore, the research community refers to these large-sized PLMs as large language models (LLMs).
@@
Figure 2: An example of LLM-based query rewriter for ad-hoc search. The example is cited from the Query2Doc paper [65]. LLMs are used to generate a passage to supplement the original query, where $N=0$ and $N>0$ correspond to zero-shot and few-shot scenarios.
@@ 3 Query Rewriter
Query rewriter, functioning as an essential preprocessing component for search engines, increases the accuracy of retrieval systems through the refinement of initial queries [66]. This mechanism, also known as query expansion or reformulation, holds a pivotal position in search engine operations. In the context of ad-hoc retrieval, the design of a query rewriter aims to mitigate the vocabulary mismatch problem by enriching original queries with semantically related terms. As conversational search evolves, query rewriters have evolved to interpret user intent and previous dialogues, thereby enabling context-sensitive queries. In this survey, the term “query rewriter” is used to refer to any technique that improves retrieval performance through query modification.
Traditional query rewriting strategies primarily include techniques such as utilizing lexical knowledge bases [67, 68, 69, 70, 71] and pseudo-relevance feedback [72, 73, 74]. However, these methods are limited due to the inadequate capabilities of knowledge models and the presence of noisy signals from coarse matching between the query and the top-$k$ retrieved documents. LLMs, pretrained on vast datasets, demonstrate considerable advancements in the breadth of knowledge and language understanding, positioning them as an excellent resource for query rewriting tasks. In the subsequent sections, we provide a comprehensive review of recent research that applies LLMs to query rewriting.
TABLE I: Overview of existing LLM-based query rewriting methods. “Knowledge” denotes the source of information the method employs for query rewriting. “SFT” and “RL” denotes supervised fine-tuning and reinforcement learning, respectively. $\mathcal{Q}$”, $\mathcal{K}$”, and $\mathcal{A}$” refer to question, keyword, and answer-incorporated passages, respectively.
Scenario Knowledge Approach Format Method
Ad-hoc LLMs Prompting $\mathcal{A}$ HyDE [75]
LLM Prompting $\mathcal{A}$ Jagerman et al. [76]
LLM Prompting $\mathcal{A}$ Query2Doc [65]
LLM Prompting $\mathcal{A}$ Baek et al. [77]
LLM Prompting $\mathcal{Q}$ Alaofi et al. [78]
LLM Prompting $\mathcal{K}$ Li et al. [79]
LLM RL $\mathcal{K}$ Ma et al. [80]
LLM SFT & RL $\mathcal{K}$ BEQUE [81]
LLM + Corpora Prompting $\mathcal{K}$ GRF+PRF [82]
LLM + Corpora Prompting $\mathcal{A}$ GRM [83]
LLM + Corpora Prompting $\mathcal{A}$ InteR [84]
LLM + Corpora Prompting $\mathcal{A}$ LameR [85]
LLM + Corpora Prompting $\mathcal{A}$ CSQE [86]
LLM + Corpora Prompting $\mathcal{Q}$ CAR [87]
LLM + Corpora SFT & RL $\mathcal{Q}$ RaFe [88]
Conver-
sational
LLM Prompting $\mathcal{Q}$ LLMCS [89]
LLM Prompting $\mathcal{Q}$ CONVERSER [90]
LLM Prompting $\mathcal{Q}$ Ye et al. [91]
3.1 Rewriting Scenarios
@@
Figure 3: An example of LLM-based query rewriter for conversational search (cited from LLMCS [89]) . An LLM is used to generate a query and system response based on the demonstrations and previous search context. $N=0$ and $N>0$ correspond to zero-shot and few-shot scenarios respectively.
In the realm of IR, a query rewriter is primarily designed to serve two distinct scenarios: ad-hoc retrieval and conversational search. Ad-hoc retrieval aims to bridge the semantic gap between a user’s query and the potential documents. LLMs, with their extensive inherent knowledge, have proven effective in replacing traditional lexical knowledge databases [67, 68, 69, 70, 71].
For conversational search, query rewriters aim to refine a query within a conversation’s context, transforming it into isolated queries based on historical dialogues. A crucial requirement for conversational query rewriters is to address coreference resolution. Traditional query rewriting methods, trained on limited data, have shown suboptimal performance, as conversational search sessions tend to be diverse and long-tailed [92, 93]. This is particularly the case in more complex conversational search sessions. However, LLMs, with their robust context understanding capabilities, have demonstrated significant advantages in conversational query rewriting [89, 90].
By leveraging the capabilities of LLMs, researchers have been able to generate a variety of formats for rewritten queries, such as questions [78, 80, 94, 91, 88, 89, 90], answer-incorporated passages [65, 77, 83, 84, 85, 75, 76, 87, 95], and keywords [82, 96, 81, 80]. Comprehensive details for each format are available in the following part.
3.2 Formats of Rewritten Queries
The intended format for rewritten queries can vary widely based on the specific needs and the downstream retrieval system. The ultimate goal is to improve the effectiveness of IR. Typically, the formats include questions, keywords, and answer-incorporated passages.
3.2.1 Questions
Rewriting original queries into similar form questions are a natural idea of query rewriting [78, 91, 88]. Query rewriters modify original queries to new questions to make it more precise, understandable, and aligned with the user’s actual search intent. This can involve rephrasing, expanding, or simplifying the query. Recent research [78] has demonstrated the potential of using LLMs to generate query variants. Although these variants cannot cover the full range of human-generated ones, they do produce highly similar sets of relevant documents.
3.2.2 Keywords
Keywords serve as a high-level abstraction of the concepts contained within a query. Rewriting queries into keywords proves particularly effective when the downstream retriever is a sparse retriever. With specific instructions, LLMs can produce high-quality keywords or concepts for query rewriting [80, 76, 81, 82]. For example, BEQUE [81] formulates new queries as keywords for effective product searches, and Li et al. [79] introduce a two-round query rewriting process, which first generates a set of high-quality seed keywords, then utilizes these keywords to enhance the query.
3.2.3 Answer-incorporated Passages
The semantic gap between short-form queries and long-form documents has been a persistent challenge. The advent of LLMs with their inherent question-answering capabilities has introduced a novel approach to query rewriting. This approach involves initially utilizing LLMs to generate comprehensive answers to the given queries. These detailed answers are then employed to retrieve relevant passages from the corpus, thereby effectively bridging the semantic divide between short queries and long candidate documents. The prompt employed for this mechanism is typically structured as follows: “Given a question query and its potential answer passages passages, compose a passage that provides an answer to the question” [85, 84]. This approach enables a more nuanced and contextually relevant retrieval of information, enhancing the overall effectiveness of the query rewriting process [65, 77, 83, 84, 85, 75, 76].
3.3 Approaches
The utilization of LLMs in query rewriting can be categorized into three primary methodologies: prompting, supervised fine-tuning, and reinforcement learning. The prompting approaches employ specific prompts to guide the LLM’s output, providing flexibility and interpretability. The supervised fine-tuning techniques adapt pre-trained LLMs to the specific task of query rewriting. However, the scarcity of training data for query rewriting often poses a challenge. To address this issue, reinforcement learning methods utilize feedback from downstream applications, thereby improving the performance of query rewriters. In the following section, we will introduce these three methods in detail.